OpenStudy (anonymous):

the table below represents the total number of laps jessica swam in the pool this summer. week 1 18 laps week 2 28 laps week 3 38 laps week 4 48 laps if the linear function representing the number of laps jessica swam each week is f(w)=10w+8. what does the 10 represent in this function?

OpenStudy (harsimran_hs4):

look if you see 18 = 1* 10 + 8 28 = 2 * 10 + 8 38 = 3 * 10 + 8 and so on so 10 is the number of laps she swam per week after 1st week

OpenStudy (anonymous):

the number of laps jessica swam in week 0 the week of the summer that jessica swam the number of laps jessica swam in total the number of laps jessica increased her swim by each week

OpenStudy (anonymous):

those are the options and im not sure wich one to choose @harsimran_hs4

OpenStudy (harsimran_hs4):

the last one she increased her laps by 10 each week

OpenStudy (anonymous):

yes that is the correct answer !!